Suppose: Consumption is $50 Investment is $10 Government Spending is $20 Imports are $5 Exports are $5 The value of GDP is $____. Write the answer as a number. Do not include the dollar sign or decimals.
Answer #1

GDP is calculated by adding consumption, government spending, investment and net exports.

GDP = C+G+I+NX = 50+20+10+(5-5) = 80

So, the value of GDP is $80.